OpenStudy (anonymous):

I GIVE MEDALS!!!!! PLZ HELP Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution.

OpenStudy (anonymous):

Equation #1: 8 - 3x = 5x + 16 Equation #2: 9x - 14 = 8x - 15 A.) The solution to equation #1 is larger than the solution to equation #2. B.) The solution to equation #1 is smaller than the solution to equation #2. C.) The solution to equation #1 is the same as the solution to equation #2. D.) None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

D. None of the statements above describe solutions to the equations shown

OpenStudy (anonymous):

I was thinking C because if you do the work...they both equal -1

OpenStudy (anonymous):

But they don't give you a number for X there for you cant correctly work out the problem :)

OpenStudy (anonymous):

8 - 3x = 5x + 16 8x = -8 x = -1 9x - 14 = 8x - 15 x = -1

OpenStudy (anonymous):

well if you know the answer then why ask

OpenStudy (loser66):

to confirm.

OpenStudy (anonymous):

^ exactly

OpenStudy (anonymous):

Well she was kind of rude to the person? no need to be rude. we are all human beings. they tried to help and she tried to argue i am just saying

OpenStudy (nuke_ur_friend):

i;m lost

OpenStudy (loser66):

@gabernathy9797 just misunderstanding,!! instead of "check my work" she asks for help, that 's it. Please, be generous

OpenStudy (anonymous):

Either I read the questions or answers wrong cause I am confused also

OpenStudy (loser66):

the question is SOLVE AND COMPARE THE SOLUTIONS so, first off, you have to solve both them to get the solutions then compare them. That's it